OpenStudy (anonymous):

A birthday hat in the shape of a cone has a lateral surface of 100π cm2. The slant height is 25 cm. Find the radius of the hat.

OpenStudy (cornzyblack):

Do you know the Lateral Surface Area formula ?

OpenStudy (anonymous):

No just the normal surface area

OpenStudy (cornzyblack):

Well lateral Surface area = pi * r * L r = RADIUS L = slant height making r subject of formula r = Lateral Surface area/ (pi * L) r = 100/ (Pi * 25)

OpenStudy (cornzyblack):

thats about 1.273 cm